OpenStudy (anonymous):

using geometric series to find the answer S = 45450, r = 3.6, n = 6

OpenStudy (campbell_st):

well the sum of a geometric series is \[S_{n} = \frac{a(1 - r^n)}{1 - r}\] you know S, r and n so I suppose you are solving for a so making a, the 1st term, the subject you get \[a = \frac{S_{n} \times (1 -r)}{1 - r^n}\] I'd substitute and evaluate

OpenStudy (anonymous):

Thank you.